XC Quiz 2 Show all work. Label and clearly explain your answer. This is very important. 1) You must answer in order to get full credit. 2) If you do show your work, and your answer is wrong. you can still earn a substantial amount of credit depending on how serious the error is. 3) If your answer is wrong, and you dont show your work, you will get a zero. Ruth Gehrig has received contract offers from three professional aquatic b teams. The details are below. Which contract is most valuable? Salary of $500,000 per year for 7 years, starting next year Salary of $333,333 per year for 7 years, starting next year, plus a $1 million signing bonus, received immediately $350,000 per year for 3 years, starting next year. In years 4-7, the salary increases to $700,000 per year. a. b. c. The interest rate is 2% per year.
Answer #1

rate positively .. let me know if you need any clarification..

We have to compute the present value of all these three case. One with the high present value will be accepted
i ii iii iv v=i*iv vi=ii*iv vii=iii*iv
Present value computation
Year Cash flow a Cash flow b Cash flow c PVIF @ 2% a b c
0        1,000,000     1.0000                      -   1,000,000.00                      -  
1        500,000            333,333 350000     0.9804      490,196.08      326,797.06      343,137.25
2        500,000            333,333 350000     0.9612      480,584.39      320,389.27      336,409.07
3        500,000            333,333 350000     0.9423      471,161.17      314,107.13      329,812.82
4        500,000            333,333 700000     0.9238      461,922.71      307,948.17      646,691.80
5        500,000            333,333 700000     0.9057      452,865.40      301,909.97      634,011.57
6        500,000            333,333 700000     0.8880      443,985.69      295,990.16      621,579.97
7        500,000            333,333 700000     0.8706      435,280.09      290,186.44      609,392.13
Present value 3,235,995.53 3,157,328.20 3,521,034.60
Present value
a 3,235,995.53
b 3,157,328.20
c 3,521,034.60
We can see present value of option c is high therefore it should be preferred over other options